# MuSACo: Multimodal Subject-Specific Selection and Adaptation for Expression Recognition with Co-Training

### v1

@@ 4 Results and Discussion
DANN [8]
CDAN [21]
MMD [31]
@@
CAN [17]
Sub-basedtop-k [45]
CMSDA [30]
@@
Upper Bound Fine-tuning 83.6 78.5 70.0 71.6 71.5 64.5 76.8 76.9 78.5 64.8 73.6
Table 1: Accuracy of MuSACo with baselines against multimodal UDA, and MSDA methods on BioVid adapted to unlabeled data from 10 target subjects, Sub-1, .., Sub-10. (Note that $M$: Multimodal, L: Labeled and U: Unlabeled.)
@@
DANN [8]
CDAN [21]
MMD [31]
@@
CAN [17]
Sub-basedtop-k [45]
@@
Upper Bound Fine-tuning 83.4 84.5 98.2 98.1 89.7 87.6 90.0 89.6 89.4 95.7 90.6
Table 2: Accuracy of MuSACo with baselines, multimodal UDA and MSDA methods on StressID adapted to 10 unlabeled target subjects.
@@ 4.1 Experimental Protocol
Datasets. All the methods were evaluated on two challenging multimodal ER datasets for subject-based adaptation. Biovid Heat Pain (PartA) [38] dataset consists of 87 subjects with 5 classes, corresponding to BL (no pain) and four pain intensity levels from PA-1 to PA-4. MuSACo follows the same protocol as [45] to consider 77 subjects as sources and 10 subjects as targets. StressID [4] dataset comprised 65 subjects taken in a lab-controlled environment. Due to the unavailability of certain modalities in 11 of the participants, our method considers 54 individuals, which includes both modalities (visual and physiological), from which 44 source and 10 target subjects are selected. More details are provided in the supplementary material. In both datasets, visual and physiological (EDA) modalities are considered in all of our experiments.
Implementation Detail. In addition to the visual modality, the physiological modality, specifically electrodermal activity (EDA), is incorporated for both the BioVid and StressID datasets. For visual, ResNet18 [12] is used, following the same protocol as [45]. The images are resized to $\times100$ resolution. For the physiological modality, an LSTM-based 1D-CNN network is included, consisting of two convolutional layers, one LSTM layer, and one fully connected layer. In addition, the expression head is constructed using 2-MLP layers. A batch size of 32 for the target subject, and the model is trained for 20 epochs. All the target subjects are split into train/val/test sets, and all the results are reported using a target test set. To select the confident target samples, $\tau_{pl}$ is set to 0.95. For selection of source subjects, $\tau_{ss}$ is set to 0.55 (Sec. 4.4-b).
Baseline Methods. The reported results correspond to the top-1% multimodal target accuracy across all target subjects. The evaluation begins with the Lower Bound, where a model is trained solely on labeled source subjects (77 for BioVid, 44 for StressID) and directly evaluated on the target test set without any target supervision. Results are reported for visual-only, physio-only, and fusion (vision + physio) modalities. Next, the MM-UDA (blending) setting blends all the source subjects into a single domain, then adapts to the unlabeled target subject using various UDA techniques, including DANN [8], CDAN [21], MMD [31], and MuSACo (UDA), which only use co-training to generate target pseudo-labels. Furthermore, MuSACo (MSDA) is compared against MM-MSDA approaches, including CAN [17], Sub-based MSDA, and CMSDA [30]. In all methods, a 1D-CNN-LSTM network is added for the physiological modality to ensure a fair comparison with our proposed approach. Finally, results for an Upper Bound setting are reported, where models are fine-tuned using labeled multimodal data from the target subject.
@@ 4.2 Comparison with State-of-the-Art Methods
Tab. 1 shows results on the BioVid dataset. All domain adaptation approaches outperform the lower-bound baselines (visual and both), confirming the effectiveness of multimodal domain adaptation in the context of pain estimation. Notably, the Physio-only baseline outperforms many standard UDA and MSDA methods, achieving an average accuracy of 34.5%. This reflects findings in prior work [40, 37, 2], where physiological signals are shown to be strong indicators of pain intensity, especially in subtle expression settings where visual cues are often weak or ambiguous. In contrast, the Visual-only baseline performs significantly worse (28.8%), and the naïve fusion of both modalities (without adaptation) does not exceed the performance of using physiological data alone. Among all methods, our approach achieves the highest overall performance. MuSACo (UDA) outperforms both the lower-bound and UDA baselines and is competitive with MSDA methods. Further, MuSACo (MSDA) achieves even better results, with a 9.3% gain over physio-only and 7.5% over MuSACo (UDA). It also outperforms MSDA baselines, with an average gain of 9.2% over CAN [17] and Sub-basedtop-k [45], and 7% over CMSDA [30]. This shows our method effectively exploits modality complementarity, adapting to subject-specific signals in subtle pain recognition.
@@
Figure 3: Left: t-SNE visualizations. The source-only produced indistinguishable feature class clusters. Sub-based MSDA [45] reduces some noise by separating classes to some extent. MuSACo creates more separable clusters for each class. Right: Visualization of selected and non-selected source subjects for a reference target subject, with similarity scores from visual and physiological modalities.
@@
Tab. 2 presents the performance on the StressID dataset. As expected, training without adaptation (lower bound) yields significantly lower accuracy across most subjects. Notably, fusing both modalities in the lower bound produces more stable performance compared to using either modality alone. This is due to the complementary nature of the modalities; when the visual modality underperforms, the physiological input often compensates, resulting in more balanced performance across subjects. Among existing methods, CAN achieves the highest performance on Sub-2, 8, whereas Sub-basedtop-k gains the performance on Sub-3 that is comparable with our method. MuSACo (UDA) improved the performance on 3 out of 10 subjects while, on average, attaining a similar performance as MSDA methods. MuSACo (MSDA) outperformed every other method on all target subjects, with an overall gain of 15.7% from lower-bound (both), 7.4% from MMD, 6.7% from MuSACo (UDA) and CAN, and 6.6% from Sub-based MSDA.
@@ 4.3 Visualizations
In Fig. 3 (left), t-SNE [35] shows how different methods separate class embeddings. The source-only model yields noisy, overlapping features. Subject-based MSDA [45] improves clustering but struggles with overlaps (e.g., PA3 in PA4). In contrast, MuSACo better clusters same-class samples and separates opposing ones. Fig. 3 (right) shows the closest and furthest source subjects to a target (Sub-10, BioVid), based on similarity scores from visual and physiological modalities. A source is selected if either modality exceeds the threshold $\tau_{ss}$, e.g., Sub-1/2 (visual) and Sub-3/4 (physio). Sub-5 to Sub-8 are excluded as both scores fall below $\tau_{ss}$ (see Sec. 4.4(b) for threshold selection).
@@ 4.4 Ablation Studies
(a) Impact of difference loss components. To evaluate the contribution of each loss component, an incremental ablation study is conducted by gradually introducing loss terms. As shown in Tab. 3, adding the unsupervised target loss $\mathcal{L}^{t}$ significantly improves accuracy over the source-only baseline. Further gains are achieved by incorporating the class-agnostic loss $\mathcal{L}_{agn}$ and the class-aware loss $\mathcal{L}_{aw}$, both of which contribute to improving the adaptation process in the absence of target labels. The combination of all loss terms achieves the highest accuracy of 40.7%.
(b) Impact of selection of source subjects threshold. As shown in Fig. 4, setting $\tau_{ss}=0$ disables the source selection mechanism, including all source subjects (77 for Biovid), achieving an average accuracy of 37.6%. As $\tau_{ss}$ increases, fewer but more relevant source subjects are selected based on similarity scores, improving performance. The best accuracy of 39.8% is achieved at $\tau_{ss}=0.55$, where only 21 source subjects are selected on average. This demonstrates the effectiveness of co-training in identifying complementary and informative sources for improved target adaptation. A breakdown of subject-wise selection of each target subject are provided in the suppl. material.
@@
$\mathcal{L}^{s}$ $\mathcal{L}^{t}$ $\mathcal{L}_{agn}$ $\mathcal{L}_{aw}$ Acc
32.1
36.5
37.2
38.0
40.7
Table 3: Impact of individual loss components.
@@ 5 Conclusion
Domain adaptation methods typically focus on individual modalities, or in the case of multimodal MSDA, they aggregate data from all source subjects, losing critical subject-specific cues necessary for effective adaptation to unlabeled target data. In this paper, MuSACo, a multimodal subject-specific MSDA is introduced that effectively leverages subject-level information for a personalized adaptation. In particular, MuSACo selects the most similar source subjects based on multimodal similarity and applies co-training to generate pseudo-labels for the target. Class-aware and class-agnostic losses are combined to leverage both confident and uncertain target samples effectively. Final alignment is achieved via feature concatenation across modalities. MuSACo is validated on challenging MER datasets and shows consistent gains over baselines, UDA (blending), and MSDA methods. MuSACo supports personalized modeling by adapting to each target subject through relevant sources, making it well-suited for health-related applications, where subject-specific cues are crucial. While the current method requires training a separate model per target subject, future work will explore lightweight or continual adaptation approaches to reduce retraining overhead and enable faster subject-specific adaptation.
@@ 6.2 Hyperparameters
Training of backbones. momentum=0.9, weight decay=5e-4, stochastic gradient descent (SGD) optimizer [29], learning rate=1e-4 with lr scheduler (eta_min=0.00002). Weighting parameters. In MuSACo, we give different weights to different loss functions. For the contribution of class-agnostic loss $\gamma=0.5$, and for class-aware loss $\alpha=0.1$. Disentanglement. Knife is very sensitive to hyperparameters; we have explored several parameters to make it work with the expression recognition task. The most critical parameters that is selected for our experiments are: zd_dim=1024, zc_dim=77, hidden_state=512, layers=3, nb_mixture=10, with learning-rate=0.01.
6.3 Detail on Training the Source Backbones
Training of the source backbones (visual and physiological) involves disentangling the identity-related features from the expression task. Our method is inspired by KNIFE[24], a fully differentiable entropy estimator, which we adapted for disentanglement in a multi-modal framework. The KNIFE estimator optimizes the backbones (visual and physiological) by decoupling non-task-related information through individual modality-specific gradient-based optimization. We leverage the KNIFE estimator to minimize the mutual information between modalities in eliminating identity features and enhancing the disentanglement of task-relevant features across distributions. We first estimate the marginal entropy of embedding $\mathbf{h}_{m}^{\textbf{s}}$ as,
@@ 6.3 Detail on Training the Source Backbones
By minimizing $\mathcal{L}_{m}^{d}$, the model is encouraged to decouple the information associated with $\acute{y}_{m}^{\textbf{s}}$ from the feature embeddings $h_{m}^{\textbf{s}}$. Furthermore, we introduce a fixed Identity-head $I_{m}$ head composed of two fully connected layers. It takes embeddings $\mathbf{h}_{m}^{\textbf{s}}$ without gradient backpropagation. $I_{m}$ works as a regularizer in conjunction with disentanglement loss to suppress non-discriminative identity-related information during target adaptation. The disentanglement loss decouples identity-related information, while the $I_{m}$ helps to constrain redundant (non-task-specific) features, ensuring the model focuses on learning relevant expression-related representations.
6.4 Datasets

### v2

@@ 4 Results and Discussion
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
@@
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
CMSDA [scalbert2021multi]
@@
27.3
28.4
30.0
35.7
Upper Bound Fine-tuning 83.6 78.5 70.0 71.6 71.5 64.5 76.8 76.9 78.5 64.8 73.6 69.1
Table 1: Accuracy and Average F1 score of MuSACo with baselines against multimodal UDA and MSDA methods on BioVid. Best Avg Acc and Avg F1 per section are highlighted in bold (ignoring supervised fine-tuning).
@@
DANN [ganin2016domain]
CDAN [long2018conditional]
MMD [sejdinovic2013equivalence]
@@
CAN [kang2020contrastive]
Sub-basedtop-k [zeeshan2024subject]
@@
53.2
54.1
62.2
Upper Bound Fine-tuning 83.4 84.5 98.2 98.1 89.7 87.6 90.0 89.6 89.4 95.7 90.6 80.4
Table 2: Accuracy and Average F1 score of MuSACo with baselines, multimodal UDA and MSDA methods on StressID adapted to 10 unlabeled target subjects. Best Avg Acc and Avg F1 per section are highlighted in bold.
Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg Acc Avg F1
Visual-only 33.3 86.0 47.1 70.5 73.0 62.0 38.2
Audio-only 35.3 35.4 44.1 65.4 60.3 48.1 36.1
Fusion 36.8 81.7 47.6 73.6 77.6 63.5 38.4
MMD [sejdinovic2013equivalence] 29.9 81.0 53.4 68.8 77.6 62.1 38.8
MuSACo (UDA) 35.4 86.6 46.7 73.5 79.4 64.3 38.7
Sub-basedtop-k [zeeshan2024subject] 33.4 84.5 54.2 71.4 79.5 64.6 38.9
MuSACo (MSDA) 39.7 95.0 54.2 79.3 81.2 69.9 41.2
Table 3: Performance of MuSACo against baselines, multimodal UDA, and MSDA methods on the BAH dataset.
@@ 4.1 Experimental Protocol
Datasets. MuSACo was evaluated on three challenging multimodal ER datasets for subject-based adaptation. BioVid Heat Pain (Part A) [walter2013biovid] contains 87 subjects with five pain levels; following [zeeshan2024subject], we use 77 as sources and 10 as targets. StressID [chaptoukaev2023stressid] comprises 65 participants in controlled settings. For our experimental setup, 44 source and 10 target subjects were used with both visual and physiological (EDA) modalities. Behavioural Ambivalence/Hesitancy (BAH) [gonzalez2025bah] consists of 224 participants from diverse demographics in uncontrolled, real-world settings. Our analysis utilizes 143 of these individuals as source subjects and five as targets with visual and audio modalities.
Implementation Detail. In addition to the visual modality, the physiological modality, specifically electrodermal activity (EDA), is incorporated for both the BioVid and StressID datasets. For visual, ResNet18 [he2016deep] is used, following the same protocol as [zeeshan2024subject]. Additional experiments using an alternative visual backbone are reported in the suppl material. For the physio modality, an LSTM-based 1D-CNN network, and for the audio modality, PANN with CNN14 [kong2020panns] model is used for the BAH dataset. All the target subjects are split into train/val/test sets, and all the results are reported using a target test set. For selection of source subjects, $\tau_{ss}$ is set to 0.55 (Sec. 4.4-b). Following prior work [deng2022robust, scalbert2021multi, ren2022multi], we set the initial $\tau_{pl}$ to 0.95 and gradually decrease it by $0.01$ every $N$ epochs to incorporate more samples as training stabilizes, with this choice validated as well as further datasets and implementation details, provided in the suppl material.
Baseline Methods. The reported results correspond to the top-1% multimodal target accuracy across all target subjects. The evaluation begins with the Lower Bound, where a model is trained solely on labeled source subjects and directly evaluated on the target test set without any target supervision. Results are reported for visual-only, physio-only, or audio-only, and fusion (vision+physio) or (vision+audio) modalities. Next, the MM-UDA (blending) setting blends all the source subjects into a single domain, then adapts to the unlabeled target subject using various UDA techniques, including DANN [ganin2016domain], CDAN [long2018conditional], MMD [sejdinovic2013equivalence], and MuSACo (UDA), which only use co-training to generate target pseudo-labels. Furthermore, MuSACo (MSDA) is compared against MM-MSDA approaches, including CAN [kang2020contrastive], Sub-based MSDA, and CMSDA [scalbert2021multi]. In all methods, we use the same backbones for every modality to ensure a fair comparison with our proposed approach. Finally, results for an Upper Bound are reported, where models are fine-tuned using labeled data from the target subject.
@@ 4.2 Comparison with State-of-the-Art Methods
Tab. 1 shows results on the BioVid dataset. All domain adaptation approaches outperform the lower bound (source-only) baselines (visual and both), confirming the effectiveness of multimodal domain adaptation in the context of pain estimation. Notably, the physio-only baseline outperforms many standard UDA and MSDA methods, achieving an average accuracy of 34.5%. This reflects findings in prior work [werner2017analysis, waligora2024joint, aslam2024distilling], where physio signals are shown to be strong indicators of pain intensity, especially in subtle expressions where visual cues are often weak or ambiguous. Similarly, our visual-only and fusion baselines yield weaker accuracy than physio-only. Nonetheless, fusion improves recall on harder classes, resulting in a higher Avg F1 score (28.1%) compared to physio-only (26.4%), despite its slightly lower accuracy. Among all methods, our approach achieves the highest overall performance. MuSACo (UDA) outperforms both the lower-bound and UDA baselines and is competitive with MSDA methods. Further, MuSACo (MSDA) achieves even better results, with a 9.3% gain over physio-only and 7.5% over MuSACo (UDA). It also outperforms MSDA baselines, with an average gain of 9.2% over CAN [kang2020contrastive] and Sub-basedtop-k [zeeshan2024subject], and 7% over CMSDA [scalbert2021multi]. This shows our method effectively exploits modality complementarity, adapting to subject-specific signals in subtle pain recognition.
@@
Figure 3: Left: t-SNE visualizations. The source-only produced indistinguishable feature class clusters. Sub-based MSDA [zeeshan2024subject] reduces some noise by separating classes to some extent. MuSACo creates more separable clusters for each class. Right: Visualization of selected and non-selected source subjects for a reference target subject, with similarity scores from visual and physiological modalities.
@@
Tab. 2 presents the performance on the StressID dataset. As expected, training without adaptation (lower bound) yields significantly lower accuracy across most subjects. Notably, fusing both modalities in the lower bound produces more stable performance compared to using either modality alone. This is due to the complementary nature of the modalities; when the visual modality underperforms, the physio often compensates, resulting in more balanced performance across subjects. Among existing methods, CAN achieves the highest performance on Sub-2, 8, whereas Sub-basedtop-k gains the performance on Sub-3 that is comparable with our method. MuSACo (UDA) improved the performance on 3 out of 10 subjects while, on average, attaining a similar performance as MSDA methods. MuSACo (MSDA) outperformed every other method on all target subjects, with an overall gain of 15.7% from lower-bound (both), 7.4% from MMD, 6.7% from MuSACo (UDA) and CAN, and 6.6% from Sub-based MSDA. Tab. 3 reports results on the BAH dataset. It can be seen that the visual modality is much stronger than audio, but fusion leverages complementary information across both modalities to achieve slight gains in performance over unimodal baselines. Among the UDA methods, MMD improves over the lower bound, while MuSACo (UDA) further stabilizes adaptation across subjects. In the MSDA setting, Sub-based demonstrates competitive results; however, MuSACo (MSDA) consistently outperforms all baselines, surpassing the lower-bound fusion by 6.4% in average accuracy and 2.3% in F1 score. Subject-wise F1 scores were provided in suppl. material.
@@ 4.3 Visualizations
In Fig. 3 (left), t-SNE [van2008visualizing] shows how different methods separate class embeddings. The source-only model yields noisy, overlapping features. Subject-based MSDA [zeeshan2024subject] improves clustering but struggles with overlaps (e.g., PA3 in PA4). In contrast, MuSACo better clusters same-class samples and separates opposing ones. Fig. 3 (right) shows the closest and furthest source subjects to a target (Sub-10, BioVid), based on similarity scores from visual and physiological modalities. A source is selected if either modality exceeds the threshold $\tau_{ss}$, e.g., Sub-1/2 (visual) and Sub-3/4 (physio). Sub-5 to Sub-8 are excluded as both scores fall below $\tau_{ss}$ (see Sec. 4.4(b) for threshold selection).
@@ 4.4 Ablation Studies
(a) Impact of difference loss components. An incremental ablation study is conducted by gradually introducing loss terms under the same configuration and weighting hyperparameters as the final model. Tab. 4 shows that adding the target loss $\mathcal{L}^{t}_{unsup}$ significantly improves Accuracy and F1 scores over the source-only baseline. Further gains are achieved by incorporating the class-aware loss $\mathcal{L}_{aw}$ and class-agnostic loss $\mathcal{L}_{agn}$, both of which contribute to improving the adaptation process in the absence of target labels. The combination of all loss terms achieves the highest performance, reaching 43.8% Accuracy and 35.7% F1. (b) Impact of selection of source subjects threshold. In Fig. 4, setting $\tau_{ss}$=0 disables the selection mechanism, including all source subjects (77 for Biovid), achieving an average accuracy of 37.6%. As $\tau_{ss}$ increases, fewer but more relevant source subjects are selected based on similarity scores, improving performance. The best accuracy of 39.8% is achieved at $\tau_{ss}$=0.55, where only 21 source subjects are selected on average. This shows the effectiveness of co-training in identifying complementary informative sources for better target adaptation. Subject-wise results are provided in the suppl. material.
@@
$\mathcal{L}^{s}$ $\mathcal{L}^{t}_{unsup}$ $\mathcal{L}_{aw}$ $\mathcal{L}_{agn}$ Avg Acc Avg F1
32.5 28.1
36.6 29.2
38.1 29.7
39.3 31.2
43.8 35.7
Table 4: Impact of individual loss components. Average performance across all subjects of the BioVid dataset.
@@ 5 Conclusion
Domain adaptation methods typically focus on individual modalities, or in the case of multimodal MSDA, they aggregate data from all source subjects, losing critical subject-specific cues necessary for effective adaptation to unlabeled target data. In this paper, MuSACo, a multimodal subject-specific MSDA is introduced that effectively leverages subject-level information for a personalized adaptation. In particular, MuSACo selects the most similar source subjects based on multimodal similarity and applies co-training to generate PLs for the target. Class-aware and class-agnostic losses are combined to leverage both confident and uncertain target samples effectively. Final alignment is achieved via feature concatenation across modalities. MuSACo is validated on challenging MER datasets, showing consistent gains across methods. MuSACo supports personalized modeling by adapting to each target subject through relevant sources, making it well-suited for health-related applications. Future work will explore test-time adaptation to reduce per-subject retraining and enable faster subject-specific adaptation.
Acknowledgment
This research was partially supported by the Natural Sciences and Engineering Research Council of Canada, Fonds de recherche du Québec – Santé, Canada Foundation for Innovation, and the Digital Research Alliance of Canada.
@@ 6 Supplementary
7 Additional Implementation Details
For visual, ResNet18 [he2016deep] network, physio network based on an LSTM-based 1D-CNN network, consisting of two convolutional layers, one LSTM layer, and one fully connected layer. For the audio modality, we employ the Pretrained Audio Neural Network (PANNs) [kong2020panns] framework, specifically the CNN14 architecture, pretrained on AudioSet. Additionally, the expression head is constructed using two 2-MLP layers for each network. A batch size of 32 for the target subject with images resized to $\times100$ resolution, and the model is trained for 20 epochs.
@@ 7.2 Hyperparameters
Training of backbones. momentum=0.9, weight decay=5e-4, stochastic gradient descent (SGD) optimizer [ruder2016overview], learning rate=1e-4 with lr scheduler (eta_min=0.00002). Weighting parameters. In MuSACo, we give different weights to different loss functions. For the contribution of class-agnostic loss $\gamma=0.5$, and for class-aware loss $\alpha=0.1$. Disentanglement. Knife is very sensitive to hyperparameters; we have explored several parameters to make it work with the expression recognition task. The most critical parameters that is selected for our experiments are: zd_dim=1024, zc_dim=77, hidden_state=512, layers=3, nb_mixture=10, with learning-rate=0.01.
8 Detail on Training the Source Backbones
Figure 5: Training of disentanglement with knife-loss and identity-head.
Training of the source backbones (visual and physiological) involves disentangling the identity-related features from the expression task. Our method is inspired by KNIFE[pichler2022differential], a fully differentiable entropy estimator, which we adapted for disentanglement in a multi-modal framework. The KNIFE estimator optimizes the backbones (visual and physiological) by decoupling non-task-related information through individual modality-specific gradient-based optimization. We leverage the KNIFE estimator to minimize the mutual information between modalities in eliminating identity features and enhancing the disentanglement of task-relevant features across distributions. We first estimate the marginal entropy of embedding $\mathbf{h}_{m}^{\textbf{s}}$ as,
@@ 8 Detail on Training the Source Backbones
By minimizing $\mathcal{L}_{m}^{d}$, the model is encouraged to decouple the information associated with $\acute{y}_{m}^{\textbf{s}}$ from the feature embeddings $h_{m}^{\textbf{s}}$. Furthermore, we introduce a fixed Identity-head $I_{m}$ head composed of two fully connected layers, as shown in Figure 5. It takes embeddings $\mathbf{h}_{m}^{\textbf{s}}$ without gradient backpropagation. $I_{m}$ works as a regularizer in conjunction with the disentanglement loss to suppress non-discriminative identity-related information during target adaptation. The disentanglement loss decouples identity-related information, while the $I_{m}$ helps to constrain redundant (non-task-specific) features, ensuring the model focuses on learning relevant expression-related representations.
9 Datasets
9.1 Biovid Heat Pain
[walter2013biovid] dataset consists of 87 subjects captured in a controlled environment. It consists of five classes, including: No-Pain, PA-1, PA-2, PA-3, and PA-4. Every individual recorded 20 videos per class, corresponding to 100 videos per subject. We follow the same protocol as [werner2017analysis], where it eliminates the initial 2 seconds from the video, which does not show any spike indicating pain. In our experiment, we follow [zeeshan2024subject] to consider 77 subjects as sources and 10 subjects as targets: Sub-1 (081014_w_27), Sub-2 (101609_m_36), Sub-3 (112009_w_43), Sub-4 (091809_w_43), Sub-5 (071309_w_21), Sub-6 (073114_m_25), Sub-7 (080314_w_25), Sub-8 (073109_w_28), Sub-9 (100909_w_65), Sub-10 (081609_w_40). The label distribution statistics for the test set of Biovid were shown in Tab. 5. Each subject has nearly balanced distributions across the five classes (neutral + 4 pain levels), confirming the dataset is overall class-balanced.
9.2 StressID